OpenStudy (anonymous):

Define independent and mutually exclusive events mathematically by using P(A) & P(B) for events A & B

OpenStudy (sidsiddhartha):

"independent events" means they are independent of each other P(A or B) = P(A)+P(B) - P(A and B) and mutually exclusive events" means they are complements of each other, either A or B happens so P(A or B)=P(A)+P(B) and P(A and B)=0

OpenStudy (sidsiddhartha):

for independent events P(A and B)=P(A)*P(B)

OpenStudy (anonymous):

if A & B are independent then P(A/B) =P(A) that is P(A/B) =P(A & B)/P(B) [ probability of occurence of A when B ahs already occurred] P(A/B) =P(A) P(A & B)/P(B) = P(A) hence condition for independent event is P(A & B) = P(B) * P(A)